# Boltzmann-Shannon Index: A Geometric-Aware Measure of Clustering Balance

### v2

@@ IV.2 Synthetic Gaussian mixtures
We begin with three deliberately constructed two-dimensional data sets, each generated from a mixture of three isotropic Gaussians (Fig. 3). The first mixture is balanced and well-separated, with comparable cluster sizes and nearly identical variances. The second introduces a moderate size imbalance while maintaining nearly identical variances. The third combines a strong imbalance in population with substantial overlap, yielding a visibly degenerate and challenging partition.
@@
In each case, we applied standard $K$-Means clustering with $K=3$, using $k$-means initialization and multiple restarts (empty-cluster action = singleton). The resulting labels were then passed directly to the Boltzmann–Shannon Index, which employs the SVD-based geometric measure $q$ described in Section III.
@@
This progression demonstrates that, without any dataset-specific tuning, the BSI reliably captures the degree to which cluster occupancy is aligned with the geometric spread each cluster occupies in the feature space, which is an intuitively desirable property that cannot be recovered by frequency-only validity measures.
@@
Figure 3: Synthetic 2D Gaussian mixtures illustrating the behavior of the Boltzmann–Shannon Index. A: balanced and well-separated clusters (high BSI). B: moderately imbalanced sizes (intermediate BSI). C: strongly imbalanced and overlapping clusters (low BSI).
IV.3 Iris benchmark dataset
@@
In this subsection, we apply the Boltzmann–Shannon Index to the classic Iris dataset fisher1936use, which contains 150150 observations with four continuous features describing three species–––setosa, versicolor, virginica–––each represented by exactly 5050 samples.
@@
Although the ground-truth species labels are known, we deliberately ignore them during clustering and performed standard $K$-means clustering with $K=3$, using 200 independent restarts and the singleton remedy for empty clusters. The best run, which was selected post-hoc as the one maximizing the BSI, recovered a partitioning that perfectly separates setosa and merges only a few versicolor-virginica points, yielding the values reported in Table 1.
@@
Boltzmann–Shannon Index (proposed) 0.9901
Boltzmann–Shannon Index (ground-truth labels) 0.9952
@@
Calinski–Harabasz index 561.6
Davies–Bouldin index 0.6620
@@
The Boltzmann–Shannon Index reaches 0.990, a value extremely close to the theoretical maximum of 1.0 and only marginally below the value of 0.9952 obtained when using the true species labels. This near-unity score reflects the fact that the recovered clusters are not only of almost perfectly equal cardinality but also occupy nearly identical effective volumes in the four-dimensional feature space, as measured by the SVD-based geometric distribution $q$.
@@
In contrast, the Boltzmann–Shannon Index simultaneously verifies that (i) the clusters are evenly populated and (ii) each cluster occupies a comparable volume in the feature space. Because the Iris dataset exhibits this ideal alignment between frequency and geometry, the BSI approaches unity. Crucially, even slight perturbations that introduce either population imbalance or geometric distortion, such as compressing one cluster or spreading another, would cause a noticeable drop in BSI well before traditional metrics register significant change. This sensitivity confirms that the proposed index is a stricter and more complete diagnostic of “balanced and geometrically meaningful” partitions than the classical alternatives.
IV.4 Resource-allocation fairness under extreme population imbalance
@@
Similarly to what we did before, we introduce a single fairness parameter $\beta\in[-1,1]$ that continuously interpolates between perfect proportionality and perfect inversion. Define the resource distribution $r(\beta)$ (which will represent the geometric probability $q$) as the convex combination
$q=r(\beta)$ $=$ $\frac{1+\beta}{2}\,p_{\text{pop}}+\frac{1-\beta}{2}\,p_{\text{rev}}$ (6)
$=$ $\frac{1+\beta}{2}\begin{pmatrix}0.950\\ 0.049\\ 0.001\end{pmatrix}+\frac{1-\beta}{2}\begin{pmatrix}0.001\\ 0.049\\ 0.950\end{pmatrix}$,
@@
where $p_{\text{rev}}$ simply reverses the population order. Consequently, $\beta=+1$ allocates resources exactly proportionally to population (ideal fairness), $\beta=0$ yields uncorrelated random allocation, and $\beta=-1$ gives almost the entire resource pool (99.9 %) to the tiniest community of 0.1% of the population (perfect injustice).
We generate N=50  [^0] N=50\,000 synthetic points whose two-dimensional coordinates reflect both population membership and per-capita resource level (so that the SVD-based geometric measure $q$ accurately captures the intended resource spread), assign the fixed labels $\{1,2,3\}$ according to $p_{\text{pop}}$, and evaluate the Boltzmann–Shannon Index as a function of $\beta$.
@@
The resulting curve (Fig. 4) reveals a smooth, monotonic, and highly informative relationship between the fairness parameter $\beta$ and the Boltzmann–Shannon Index. At complete inversion ($\beta=-1$), when nearly all resources are allocated to the tiniest community ($0.1\%$ of the population), the index collapses to BSI $\leq 0.06$. As $\beta$ increases toward neutrality, the BSI gradually increases to approximately 0.70 at $\beta=0$ (uncorrelated allocation), and, eventually, reaches BSI $\geq 0.98$ only under strict proportionality ($\beta=+1$), reflecting the near-perfect alignment between demographic weight and geometric spread. The absence of abrupt transitions is itself instructive: in the presence of extreme demographic skew (95 %–4.9 %–0.1 %), truly catastrophic injustice is required to drive the index to near-zero, yet even modest departures from perfect proportionality are consistently and visibly penalized, producing a smooth but uncompromising gradient of fairness.
@@
This smooth yet unforgiving behavior confers two immediate practical advantages for quantifying justice in resource allocation. First, the index produces a continuous and interpretable scalar value in $[0,1]$ that can be reported alongside established measures of inequality, such as the Gini coefficient gini1912variabilita, Theil index theil1967economics, and Atkinson index atkinson1970measurement. At the same time, it captures a complementary notion of fairness grounded jointly in frequency (share of resources or beneficiaries) and geometry (distribution over the underlying allocation space). Moreover, unlike purely frequency-based metrics, the BSI accounts simultaneously for both demographic weight and the effective “living space” (or opportunity space) available to each group, making it particularly sensitive to scenarios in which a small elite occupies a disproportionately large region of the state space while the majority is compressed into a narrow, resource-poor domain. Second, and more importantly, the Boltzmann–Shannon Index is differentiable with respect to the allocation vector $r$, since both the frequency distribution $p$ and the SVD-based geometric distribution $q$ vary smoothly with $r$. This property enables the BSI to be incorporated directly as a regularization term in constrained optimization formulations for resource allocation. Such formulations naturally arise in public-policy design (health-care budgets, educational funding, carbon-emission allowances), disaster-relief distribution, humanitarian aid allocation, and machine-learning fairness constraints (e.g., demographic parity with geometric awareness). Because maximizing the BSI is equivalent to minimizing the Jensen–Shannon divergence between demographic and geometric distributions, the resulting optima favour allocations in which each community receives resources in proportion to both its population and its effective spread in outcome space. This yields a principled notion of equity that goes well beyond simple per-capita uniformity, rewarding allocations that balance demographic weight with the breadth of opportunities or needs reflected in the geometry of the outcome space.
Thus, in sharply skewed populations, the Boltzmann–Shannon Index does not merely diagnose imbalance, but it also provides an operationally meaningful, gradient-friendly objective that can be directly embedded into the optimization machinery of modern policy and algorithmic governance.
@@
Figure 4: Boltzmann–Shannon Index as a function of the fairness parameter $\beta$ for population shares 95.0 %–4.9 %–0.1 %. The index is maximal ($\approx 0.98$) only under strictly proportional allocation ($\beta=+1$) and collapses to nearly zero when resources are inverted toward the smallest community ($\beta=-1$).
V Conclusion
@@
In this paper, we introduced the Boltzmann–Shannon Index as a simple but effective way to quantify how well a clustering reflects both the frequency and the geometry of data in continuous spaces. By pairing the empirical distribution of cluster labels with a geometric distribution derived from SVD-based volume estimates, the index provides a direct measure of how evenly the state space is partitioned. When these two distributions agree, the index approaches its ideal value (i.e., 1), while when they diverge, the index drops smoothly and predictably towards the lowest possible value (i.e., 0). This behavior gives an immediate sense of whether a partition is “balanced” in a meaningful geometric sense, something that traditional clustering indices are not designed to detect.
@@
Finally, the differentiability of the BSI with respect to the underlying allocations suggests a practical role beyond diagnostic use. It can be incorporated directly into optimization frameworks where fairness or balanced representation is a design objective, providing a smooth penalty that favors partitions aligned with both demographic weight and geometric extent. We expect this dual perspective of considering frequency and geometry together to be increasingly valuable in applications involving mixed continuous data, coarse-grained dynamical systems, and modern resource-allocation problems.
Looking ahead, several natural extensions emerge from this work. One important direction is the treatment of mixed continuous–discrete data, where geometric structure coexists with categorical attributes; adapting the BSI to such settings would broaden its applicability across many real-world domains, including important machine learning applications. A second direction concerns domain-specific applications, including coarse-grained dynamical systems, high-resolution imaging, and policy-driven resource allocation, where geometry-aware fairness or interpretability constraints may play an operational role.

### v3

@@ III.2 Synthetic Gaussian mixtures
Fig. 3 includes three deliberately constructed two-dimensional data sets (a)-(c), each generated from a mixture of three isotropic Gaussians. The first mixture (a) is balanced and well-separated, with comparable cluster sizes and nearly identical variances; the second (b) introduces a moderate size imbalance while maintaining nearly identical variances; and the third (c) combines a strong imbalance in population with substantial overlap, yielding a visibly degenerate and challenging partition.
@@
In each case, we applied standard $K$-Means clustering with $K=3$, using random initialization and multiple restarts (empty-cluster action = singleton). The resulting labels were then passed directly to the Boltzmann–Shannon Index, which employs the SVD-based geometric measure q described in Sec. II.
@@
This example shows that the BSI reliably captures the degree to which cluster membership is aligned with the geometric spread each cluster occupies in the feature space, which is an intuitively desirable property that cannot be recovered by frequency-only validity measures.
Refer to caption Refer to caption
(a) (b)
@@
Figure 3: Synthetic 2D Gaussian mixtures illustrating the behavior of the Boltzmann–Shannon Index. (a) balanced and well-separated clusters (high BSI); (b) moderately imbalanced sizes (intermediate BSI); and (c) strongly imbalanced and overlapping clusters (low BSI).
III.3 Iris benchmark dataset
@@
We also apply the Boltzmann-Shannon Index to the classic Iris dataset fisher1936use, which contains $150$ observations with four continuous-valued features describing the three species of setosa, versicolor, and virginica, each represented by exactly $50$ samples, an objectively near equal density case.
@@
Although the ground-truth species labels are known, we deliberately ignore them during clustering and performed standard $K$-means clustering with $K=3$, using $200$ independent restarts and the singleton remedy for empty clusters. The best run, i.e. the one maximizing the BSI, resulted in a partition that perfectly separates setosa and merges only a few versicolor-virginica points, yielding the values reported in Table 1.
@@
Boltzmann-Shannon Index (proposed) 0.9901
Boltzmann-Shannon Index (ground-truth labels) 0.9952
@@
Calinski-Harabasz index 561.6
Davies-Bouldin index 0.6620
@@
The Boltzmann-Shannon Index applied to the best result from $K$-means reaches $0.990$, a value extremely close to the theoretical maximum of $1.0$ and only marginally below the value of $0.9952$ that is obtained when using the true species labels. This near-unity score reflects the fact that the recovered clusters are not only of almost perfectly equal cardinality but also occupy nearly identical effective volumes in the four-dimensional feature space, as measured by the SVD-based geometric distribution q.
@@
In contrast, the Boltzmann-Shannon Index simultaneously verifies that (i) the clusters are evenly populated and (ii) each cluster occupies a comparable volume in the feature space. Because the Iris dataset exhibits this ideal alignment between frequency and geometry, the BSI approaches unity. Crucially, even slight perturbations that introduce either population imbalance or geometric distortion, such as compressing one cluster or spreading another, would cause a noticeable drop in BSI well before traditional metrics register a significant change. This sensitivity confirms that the proposed index is a stricter and more complete diagnostic of “density balanced” partitions than the classical alternatives.
III.4 Resource-allocation fairness under extreme population imbalance
@@
Similar to Subsec. III.1, we introduce a single fairness parameter $\beta\in[-1,1]$ that continuously interpolates between perfect proportionality and perfect inversion by defining a resource distribution $r(\beta)$ (which will be used to generate geometric probability distributions q for each value of $\beta$) as the convex combination
$r(\beta)=$ $\frac{1+\beta}{2}\,\textbf{p}_{\text{pop}}+\frac{1-\beta}{2}\,\textbf{p}_{\text{rev}}$
$=$ $\frac{1+\beta}{2}\begin{pmatrix}0.950\\ 0.049\\ 0.001\end{pmatrix}+\frac{1-\beta}{2}\begin{pmatrix}0.001\\ 0.049\\ 0.950\end{pmatrix}$, (6)
@@
where $\textbf{p}_{\text{rev}}$ simply reverses the population order. Consequently, $\beta=+1$ represents resources being allocated exactly proportionally to population (total fairness); $\beta=0$ yields uncorrelated random allocation; and $\beta=-1$ gives almost the entire resource pool ($95\%$) to the tiniest community of $0.1\%$ of the population (extreme inequality).
For a range of $\beta$ values, we generate $N=500,000$ points (two-dimensions) sampled from a Gaussian distribution, with coordinates representing the resource allocation parameters. Each dimension of the resource space corresponds to a distinct type of resource. The label for each point is assigned according to the population probability distribution $p_{\text{pop }}$.
To ensure that the geometric structure of the resources scales appropriately in accordance with Eq. 6, the allocation vector $r$ is computed as a function of $\beta$. For each $\beta$ value, the coordinates of the points are rescaled as $X_{i}=U_{i}R_{i}V_{i}^{T}$, where $U_{i}$ and $V_{i}$ are unitary matrices obtained from the singular value decomposition (SVD) of the group (community) matrix $X_{i}$, and $R_{i}$ is a diagonal matrix with the allocation parameters $r_{i}(\beta)$ (corresponding to the $\mathrm{i}^{th}$ group) along the diagonal.
The resulting relationship (see Figure 4) exhibits a smooth and highly informative curve linking the fairness parameter $\beta$ to the Boltzmann-Shannon Index.
@@
At complete inversion ($\beta=-1$), when nearly all resources are allocated to the tiniest community ($0.1\%$ of the population), the index collapses to $\operatorname{BSI}(-1)\approx 0.06$. As $\beta$ increases toward neutrality, the BSI gradually increases to approximately $0.70$ at $\beta=0$ (an interesting value for uncorrelated allocation), and, eventually, reaches $\operatorname{BSI}(1)\approx 0.98$ only under strict proportionality, reflecting the near-perfect alignment between demographic weight and geometric spread. The absence of abrupt transitions is itself instructive: in the presence of extreme demographic skew $(95\%,4.9\%,0.1\%)$, truly catastrophic inequality is required to drive the index to near-zero, yet even modest departures from perfect proportionality are consistently and visibly penalized, producing a smooth but uncompromising gradient of fairness. Of particular interest is the relatively more positive (greater than $0.5$) value of uncorrelated resource allocation.
@@
This smooth behavior confers two immediate practical advantages for quantifying inequality in resource allocation. First, the index produces a continuous and interpretable scalar value in $[0,1]$ that can be reported alongside other established measures, such as the Gini coefficient gini1912variabilita, Theil index theil1967economics, and Atkinson index atkinson1970measurement. At the same time, it captures a complementary notion of fairness grounded jointly in frequency (the number of members sharing the portion of resources) and geometry (the relative distribution of the underlying allocated resources). Moreover, unlike purely frequency-based metrics, the BSI accounts simultaneously for both demographic weight and the effective opportunity space available to each group, making it particularly sensitive to scenarios in which a small elite occupies a disproportionately large region of the state space while the majority is compressed into a narrow, resource-poor domain.
Second, and more importantly, the Boltzmann–Shannon Index is effectively differentiable with respect to the allocation vector $r$, since both the frequency distribution p and the SVD-based geometric distribution q vary smoothly with $r$ (at least numerically). This property enables the BSI to be incorporated directly as a regularization term in constrained optimization formulations for resource allocation. Such formulations naturally arise in public-policy design (health-care budgets, education funding), disaster-relief distribution, and humanitarian aid allocation. Because maximizing the BSI is equivalent to minimizing the Jensen–Shannon divergence between demographic and resource distributions, the resulting optima favor allocations in which each community receives resources in proportion to both its population and its effective spread in outcome space. Although this may not always be the goal of optimization, it can provide a useful regularization term.
In addition, maximizing the Boltzmann–Shannon Index is not universally desirable. In certain domains, a deliberately low BSI may be preferable. A prominent example is the analysis of carbon-emissions across entities (countries, firms, or individuals). In this context, the optimal outcome may be one in which the overwhelming majority of emitters produce very little carbon, and a small number of high-emission activities are those that can be managed or regulated. Thus, depending on the normative goal, e.g., equity in opportunity versus concentration of harmful outcomes, the BSI can be either maximized (as a fairness regularizer) or minimized (as a harm reduction regularizer).
@@
Figure 4: Boltzmann–Shannon Index as a function of the fairness parameter $\beta$ for population shares 95.0 %,4.9 %,0.1 %. The index is maximal ($\approx 0.98$) under strictly proportional allocations ($\beta=+1$) and collapses to nearly zero when resources are inverted toward the smallest community ($\beta=-1$).
IV Conclusion
@@
In this paper, we introduced the Boltzmann–Shannon Index as a simple but effective way to quantify how well a clustering reflects both the frequency and the geometry of data in continuous spaces. By pairing the distribution of frequencies in cluster labels with a geometric distribution derived from SVD-based volume estimates, the index provides a direct measure of how evenly the state space is partitioned. When these two distributions agree, the index approaches its maximum value (i.e., 1), while when they diverge, the index drops smoothly and predictably towards the lowest possible value (i.e., 0). This behavior gives an immediate sense of whether a partition is “density-balanced” in a meaningful geometric sense, something that traditional clustering indices are not designed to detect.
@@
Finally, the numerical differentiability of the BSI suggests a practical role beyond diagnostic use. For instance, it can be incorporated directly into optimization frameworks where fairness or balanced representation is a design objective, providing a smooth penalty that favors partitions aligned with both demographic weight and geometric extent. We expect this dual perspective of considering frequency and geometry together to be increasingly valuable in applications involving mixed data, coarse-grained dynamical systems, and modern resource-allocation problems.
Data Availability
Data and relevant code for this research work are freely available at: https://github.com/almomaa/Boltzmann_Shannon_Index